# Revela: Dense Retriever Learning via Language Modeling

### v1

@@ B.1 Analysis on the Size of LMs
AppsRetrieval 5.2 5.9 6.3 15.8 17.9 19.0 12.7 19.6 16.8
CosQA 22.8 22.6 24.2 25.7 25.5 27.6 26.5 26.8 27.7
SyntheticText2SQL 43.2 44.8 44.7 54.4 54.1 53.3 49.7 53.2 53.4
COIRCodeSearchNetRetrieval 36.8 40.5 43.2 47.9 48.1 52.6 53.7 53.7 60.1
CodeSearchNetCCRetrieval 52.2 56.7 58.9 57.6 61.2 63.6 59.6 63.2 67.0
CodeTransOceanContest 60.1 60.1 60.6 68.1 69.0 71.3 67.1 72.0 74.4
CodeTransOceanDL 33.3 33.4 32.9 32.8 33.7 34.1 32.8 32.4 33.7
StackOverflowQA 68.1 66.3 68.3 79.1 77.8 81.0 80.2 81.7 83.8
CodeFeedbackST 63.2 63.0 65.0 72.7 72.5 73.8 72.7 74.6 74.5
CodeFeedbackMT 42.4 42.1 43.1 51.1 51.8 52.3 51.2 54.4 54.6
Mean 42.7 43.5 44.7 50.5 51.2 52.9 50.6 53.2 54.6
From Table 7, we can observe that larger rank in LoRA will lead to better performance for Revela training. This demonstrates the potential scalability of Revela, when we enlarge the rank or conduct full-parameter tuning.
Table 7: Retrieval performance (NDCG@10, %) across BEIR benchmarks, using three model sizes (columns in blocks) and different LoRA ranks (r=16, 64, 256). Bold in each row indicates the highest performance for the specific task and that model size.
Revela-135M Revela-500M Revela-1B
Dataset r=16 r=64 r=256 r=16 r=64 r=256 r=16 r=64 r=256
ArguAna 38.1 38.8 39.4 41.1 40.5 41.4 44.0 44.4 45.1
ClimateFEVER 15.3 15.8 15.6 14.2 13.7 13.5 14.0 13.9 15.6
DBPedia 17.7 17.3 18.3 22.9 19.7 21.4 22.3 22.4 27.6
FEVER 33.1 35.3 35.1 49.2 47.4 50.9 54.6 59.2 60.4
FiQA2018 19.3 18.7 19.1 27.0 25.6 26.6 27.2 28.9 29.6
HotpotQA 38.5 35.6 38.5 51.6 48.0 50.6 52.7 52.6 56.0
NFCorpus 22.9 24.1 23.5 27.8 26.5 26.5 28.1 29.1 27.2
NQ 19.5 19.8 21.2 29.2 25.5 29.8 27.8 28.4 33.9
QuoraRetrieval 81.3 81.9 81.0 83.9 81.5 83.1 82.6 81.8 83.5
SCIDOCS 11.5 11.0 12.1 15.8 15.5 14.8 15.3 16.1 16.4
SciFact 55.9 57.5 57.9 64.4 64.8 65.3 68.2 70.8 71.8
TRECCOVID 57.6 61.2 60.2 61.0 57.4 58.4 59.1 61.9 60.0
Touche2020 14.6 14.1 14.4 20.9 19.2 21.1 18.8 21.5 23.3
Mean 32.7 33.1 33.6 39.2 37.3 38.7 39.6 40.8 42.3
B.2 Analysis on Batch Construction
Table 8 and 9 present the performance of Revela with different setups of training dataset construction on BEIR and CoIR, respectively.
Table 8: Retrieval performance (NDCG@10, %) across BEIR, for three model sizes (135M, 500M, 1B) and different training data setups (vanilla, random, dual).
Revela-135M Revela-500M Revela-1B
Dataset vanilla random dual vanilla random dual vanilla random dual
ArguAna 39.4 13.9 41.6 41.4 10.7 44.4 45.1 15.4 44.7
ClimateFEVER 15.6 6.3 16.3 13.5 7.9 15.5 15.6 6.6 15.1
DBPedia 18.3 1.2 18.7 21.4 2.0 21.7 27.6 3.2 25.4
FEVER 35.1 1.0 36.6 50.9 0.8 44.7 60.4 2.3 56.6
FiQA2018 19.1 0.9 21.0 26.6 2.3 27.9 29.6 1.6 34.0
HotpotQA 38.5 1.1 36.9 50.6 1.4 50.3 56.0 1.4 59.1
NFCorpus 23.5 2.6 24.7 26.5 3.7 28.4 27.2 4.3 28.7
NQ 21.2 0.5 19.7 29.8 0.6 27.4 33.9 2.0 34.5
QuoraRetrieval 80.9 0.5 82.2 83.1 0.4 85.7 83.5 0.5 85.2
SCIDOCS 12.1 1.3 11.6 14.8 2.5 15.7 16.3 2.2 16.1
SciFact 57.9 11.0 58.9 65.2 7.6 62.8 71.8 7.5 69.3
TRECCOVID 60.2 7.6 62.0 58.4 13.2 62.6 59.8 8.2 66.5
Touche2020 14.4 2.8 17.2 21.0 9.3 18.5 23.3 3.6 23.0
Mean 33.6 3.9 34.4 38.7 4.8 38.9 42.3 4.5 42.9
Table 9: Retrieval performance (%) across CoIR, for three model sizes (135M, 500M, 1B) and different training data setups (vanilla, random, dual).
Revela-135M Revela-500M Revela-1B
Dataset chunk random dual chunk random dual chunk random dual
AppsRetrieval 6.3 0.0 6.3 19.0 0.0 17.1 16.8 0.4 17.7
CosQA 24.2 0.0 23.7 27.6 0.0 27.0 27.7 5.2 25.3
SyntheticText2SQL 44.7 0.1 45.9 53.3 0.0 53.4 53.4 1.5 55.2
COIRCodeSearchNetRetrieval 43.2 0.0 41.3 52.6 0.0 51.6 60.1 0.9 58.3
CodeSearchNetCCRetrieval 58.9 0.0 59.3 63.6 0.0 62.7 66.9 1.2 66.6
CodeTransOceanContest 60.6 1.0 57.3 71.3 0.5 70.7 74.4 3.8 74.1
CodeTransOceanDL 32.9 0.7 34.2 34.1 1.2 33.8 33.7 8.3 34.0
StackOverflowQA 68.3 0.3 66.8 81.0 0.0 80.0 83.8 19.7 83.3
CodeFeedbackST 65.0 0.0 63.0 73.8 0.0 73.5 74.5 4.4 74.5
CodeFeedbackMT 43.1 0.1 41.9 52.3 0.0 52.0 54.6 2.7 55.3
Mean 44.7 0.2 44.0 52.9 0.2 52.2 54.6 4.8 54.4
B.3 Analysis on Batch sizes
Table 10 and 11 present the performance of Revela with different batch sizes on BEIR and CoIR, respectively. Although models trained on a batch size of 16 achieve higher mean performance, they cannot consistently outperform those trained with smaller batch sizes across all tasks. This finding demonstrates the possibility to adopt smaller batch sizes for Revela training.
Table 10: Retrieval performance (NDCG@10, %) across the 13 BEIR tasks for three encoder sizes (135M, 500M, 1B) and three batch sizes (bs8, bs12, bs16).
Revela-135M Revela-500M Revela-1B
Dataset bs8 bs12 bs16 bs8 bs12 bs16 bs8 bs12 bs16
ArguAna 39.7 41.1 39.4 42.9 39.9 41.4 44.4 44.4 45.1
ClimateFEVER 15.7 15.9 15.6 15.2 12.5 13.5 16.0 13.3 15.6
DBPedia 17.1 17.0 18.3 21.5 18.1 21.4 25.9 23.7 27.6
FEVER 36.0 37.0 35.1 54.2 44.4 50.9 55.4 58.7 60.4
FiQA2018 20.1 18.7 19.1 26.2 25.5 26.6 29.7 29.0 29.6
HotpotQA 36.8 35.6 38.5 50.9 46.3 50.6 55.4 53.9 56.0
NFCorpus 22.7 24.0 23.5 24.1 25.7 26.5 30.1 29.6 27.2
NQ 19.4 19.5 21.2 24.7 22.4 29.8 30.8 30.5 33.9
QuoraRetrieval 80.0 80.2 81.0 82.6 82.2 83.1 82.1 83.4 83.5
SCIDOCS 11.5 11.3 12.1 14.8 14.3 14.8 16.5 15.7 16.3
SciFact 56.9 58.0 57.9 63.2 62.2 65.2 68.8 69.3 71.8
TRECCOVID 61.0 59.7 60.2 57.3 58.5 58.4 56.7 59.8 59.8
Touche2020 16.3 14.1 14.4 20.0 20.7 21.0 24.9 22.9 23.3
Mean 33.3 33.2 33.6 38.3 36.4 38.7 41.3 41.1 42.3
Table 11: Retrieval performance (%, NDCG@10) across the 10 CoIR tasks. We report three encoder sizes (135M, 500M, 1B) and three batch sizes (bs8, bs12, bs16).
Revela-135M Revela-500M Revela-1B
Dataset bs8 bs12 bs16 bs8 bs12 bs16 bs8 bs12 bs16
AppsRetrieval 6.8 6.5 6.3 17.8 16.5 19.0 17.9 19.0 16.8
CosQA 22.2 24.0 24.2 26.1 28.9 27.6 27.9 28.6 27.7
SyntheticText2SQL 42.0 41.9 44.7 52.8 55.2 53.3 54.8 55.5 53.4
COIRCodeSearchNetRetrieval 43.5 43.4 43.2 52.5 50.8 52.6 58.8 56.2 60.1
CodeSearchNetCCRetrieval 57.9 59.8 58.9 62.0 63.3 63.6 66.2 65.5 67.0
CodeTransOceanContest 58.8 58.7 60.6 66.8 70.1 71.3 71.8 74.2 74.4
CodeTransOceanDL 34.7 32.6 32.9 33.5 33.9 34.1 32.9 33.3 33.7
StackOverflowQA 67.4 66.7 68.3 78.2 80.9 81.0 84.8 85.0 83.8
CodeFeedbackST 62.5 63.0 65.0 73.3 74.7 73.8 75.0 75.1 74.5
CodeFeedbackMT 41.6 41.7 43.1 51.8 51.6 52.3 55.1 54.1 54.6
Mean 43.7 43.8 44.7 51.5 52.6 52.9 54.5 54.7 54.6
B.4 Analysis of V-normalization
Table 12 compares Revela with and without the proposed V-normalization. We observe that across various datasets within BEIR, including V-normalization mostly outperforms its absence, highlighting the effectiveness of the design.
Table 12: NDCG@10 (%) on BEIR across three models in different sizes. Bold numbers in each row indicate the better performance between w/o v-norm vs. w/ v-norm for a given model size.
Revela-135M Revela-500M Revela-1B
Dataset w/o v-norm w/ v-norm w/o v-norm w/ v-norm w/o v-norm w/ v-norm
ArguAna 42.5 39.4 41.5 41.4 45.6 45.1
C-FEVER 14.5 15.6 12.0 13.5 12.7 15.6
DBPedia 16.1 18.3 17.1 21.4 22.8 27.6
FEVER 29.0 35.1 41.3 50.9 46.0 60.4
FiQA2018 18.2 19.1 25.1 26.6 26.9 29.6
HotpotQA 35.8 38.5 44.8 50.6 49.7 56.0
NFCorpus 23.7 23.5 26.8 26.5 29.1 27.2
NQ 16.2 21.2 19.1 29.8 22.1 33.9
QuoraRetrieval 81.5 80.9 80.4 83.1 83.3 83.5
SCIDOCS 10.7 12.1 14.6 14.8 16.5 16.3
SciFact 58.2 57.9 62.6 65.2 71.6 71.8
TRECCOVID 47.6 60.2 50.4 58.4 48.1 59.8
Touche2020 11.1 14.4 14.7 21.0 17.6 23.3
Mean 31.2 33.6 34.7 38.7 37.9 42.3

### v2

@@ C.3 Analysis on the Size of LMs
AppsRetrieval 5.9 5.2 8.2 18.2 15.0 20.5 17.9 15.7 19.4
CosQA 24.8 24.6 26.2 26.7 28.3 27.5 28.7 30.2 30.2
SyntheticText2SQL 43.9 46.2 45.7 50.8 52.0 53.7 52.1 53.0 55.0
COIRCodeSearchNetRetrieval 40.0 40.0 49.9 51.2 48.0 57.9 57.2 55.2 64.0
CodeSearchNetCCRetrieval 54.7 56.2 63.4 61.5 61.3 68.0 62.6 64.1 70.0
CodeTransOceanContest 67.0 67.6 70.9 74.0 74.0 77.6 76.8 76.8 81.1
CodeTransOceanDL 34.2 34.4 34.6 34.0 34.8 35.4 33.8 35.0 34.2
StackOverflowQA 67.1 65.4 69.2 79.5 77.0 82.5 82.2 80.9 85.7
CodeFeedbackST 64.8 64.9 63.8 73.3 73.9 74.5 75.2 75.0 76.2
CodeFeedbackMT 55.7 54.4 51.7 66.0 62.4 63.6 68.5 69.7 70.4
Mean 45.8 45.9 48.4 53.5 52.7 56.1 55.5 55.6 58.6
C.4 Mixed-domain composition
Tab. 14 and Tab. 15 report the performance of Revela trained on the mixture of the batches constructed from Wikipedia and code-related corpus, applied in § 4. We randomly sample 160,000 batches from both datasets to form the training data, and maintain all other experimental setups. Compared with Tab. 1 and Tab. 6, where Revela is trained separately in each domain, it can largely maintain the original performance when trained on the mixed-domain data. These results indicate Revela’s potential to generalize to diverse domains.
Table 14: Comparison of Revela’s performance (NDCG@10, %) across CoIR benchmark tasks for three LM sizes trained on the mixture of Wikipedia and code-related corpus. Bold indicates the best performance in each row.
Dataset Revela0.1B Revela0.5B Revela1B
AppsRetrieval 5.9 17.2 17.5
CosQA 25.8 28.0 24.5
SyntheticText2SQL 46.0 51.5 55.1
COIRCodeSearchNetRetrieval 41.1 51.5 58.3
CodeSearchNetCCRetrieval 59.3 62.7 66.6
CodeTransOceanContest 66.3 76.4 78.6
CodeTransOceanDL 34.4 35.0 34.3
StackOverflowQA 66.0 79.4 82.2
CodeFeedbackST 64.5 74.6 75.6
CodeFeedbackMT 52.8 67.7 71.2
Mean 46.2 54.4 56.4
Table 15: Comparison of Revela’s performance (NDCG@10, %) across BEIR benchmark tasks for three LM sizes trained on the mixture of Wikipedia and code-related corpus. The highest value in each row is highlighted in bold.
Dataset Revela0.1B Revela0.5B Revela1B
ArguAna 41.4 43.9 44.2
ClimateFEVER 16.2 16.1 15.3
DBPedia 18.8 21.7 25.4
FEVER 35.9 44.9 56.8
FiQA2018 21.2 28.3 34.6
HotpotQA 37.0 50.3 59.1
NFCorpus 24.5 28.7 28.7
NQ 19.7 27.4 34.5
QuoraRetrieval 82.2 85.7 85.2
SCIDOCS 11.7 15.8 16.1
SciFact 59.5 63.8 69.2
TRECCOVID 61.8 62.9 67.0
Touche2020 15.7 20.0 23.1
Mean 34.3 39.2 43.0
C.5 Out-of-domain Generalization
To further validate Revela’s generalization, we trained it on Fineweb-edu (Penedo et al., 2024), a widely-used language modeling corpus. [^7] https://huggingface.co/datasets/HuggingFaceFW/fineweb-edu Following the experimental setup from § 4.1, we trained for 320,000 batches and evaluated the models on the CoIR and BEIR benchmarks. The results in Tab. 16 and Tab. 17 show that Revela achieves competitive performance despite being trained on out-of-domain data. Notably, the Revela0.5B model scores 48.6% on CoIR, outperforming E5-PT (46.4%), which was trained on 270 million query-document pairs.
Table 16: Comparison of Revela’s performance (NDCG@10, %) across CoIR benchmark tasks for three LM sizes trained on an out-of-domain corpus, i.e., Fineweb-edu. Bold indicates the best performance in each row.
Dataset Revela0.1B Revela0.5B Revela1B
AppsRetrieval 1.9 8.6 11.3
CosQA 25.8 26.0 26.6
SyntheticText2SQL 42.0 54.3 53.7
COIRCodeSearchNetRetrieval 27.8 38.9 46.2
CodeSearchNetCCRetrieval 47.0 60.5 61.7
CodeTransOceanContest 60.2 68.6 75.2
CodeTransOceanDL 33.4 33.5 33.8
StackOverflowQA 50.6 63.1 69.0
CodeFeedbackST 52.7 70.4 70.9
CodeFeedbackMT 38.7 61.8 66.9
Mean 38.0 48.6 51.5
Table 17: Comparison of Revela’s performance (NDCG@10, %) across 13 BEIR tasks for three LM sizes trained on an out-of-domain corpus, i.e., Fineweb-edu. Bold indicates the best performance in each row.
Dataset Revela0.1B Revela0.5B Revela1B
ArguAna 36.4 36.9 38.7
ClimateFEVER 13.9 12.7 12.3
DBPedia 17.2 20.2 24.5
FEVER 30.5 45.0 53.9
FiQA2018 20.1 24.9 30.3
HotpotQA 29.7 47.5 54.2
NFCorpus 26.8 30.1 29.7
NQ 17.4 24.8 31.0
QuoraRetrieval 81.8 83.6 82.4
SCIDOCS 12.6 15.6 16.8
SciFact 59.3 65.1 68.6
TRECCOVID 65.5 66.3 63.3
Touche2020 15.8 20.9 23.0
Mean 32.9 38.0 40.7
C.6 LM performance after Revela
We evaluate the original LLaMA-3.2-1B and the model co-trained with the retriever (LLaMA-3.2-1B backboned on Wikipedia in § 4.1) on seven tasks, including ARC (Clark et al., 2018), COPA (Roemmele et al., 2011), OpenBookQA (Mihaylov et al., 2018), PIQA (Mihaylov et al., 2018), RTE (Bisk et al., 2020), WiC (Dagan et al., 2005), and Winogrande (Pilehvar & Camacho-Collados, 2019). As shown in Tab. 18, LM’s performance has been greatly preserved.
Table 18: Performance (Accuracy, %) of the LM before and after Revela on various benchmarks.
Model ARC COPA OpenBookQA PIQA RTE WiC WinoGrande Avg.
LLaMA-3.2-1B 28.0 77.0 25.0 73.0 59.5 47.5 57.5 52.5
LM-Revela 29.0 74.0 25.5 71.0 58.5 52.0 55.5 52.2
C.7 Comparison between Revelaand REPLUG
While both Revela and REPLUG involve retrievers and LMs, their core motivations differ. REPLUG distills relevance signals from the perplexity of frozen, off-the-shelf LMs. In contrast, Revela proposes a broader paradigm: to learn retrievers as an integral part of language modeling, rather than as a post hoc add-on. From a technical perspective, Revela also outperforms REPLUG regarding efficiency.
For a batch of $B$ documents (each with sequence length $L$), the cost of modeling inter-document relationships differs significantly between REPLUG and Revela:
  • REPLUG: Computes pairwise relevance between all document pairs using LM perplexity, resulting in a forward cost of $B^{2}\times(2L)^{2}$, where the pair of texts is concatenated to pass through the LMs to calculate their relevance.
  • Revela: As detailed in § 3.4, uses in-batch attention with concatenated inputs, yielding a forward cost of $B\times(2L)^{2}$. If approximately taking backward cost as the same as forward, the cost is still linearly proportional to the batch size.
Therefore, when using a reasonably large batch size (e.g., 16 as in our setup), Revela learns inter-document relevance more efficiently than REPLUG, due to its linear scaling in batch size compared to REPLUG’s quadratic cost.
Appendix D Use of LLMs
In preparing this manuscript, we employed LLM for general assistance. Its use focuses on improving the clarity and grammar of the text and helping generate the code used to create figures (e.g., matplotlib) or LaTeX tables.